# Covariant effective spacetimes of spherically symmetric electro-vacuum with a cosmological constant

### v1

@@ Abstract
An algebraic framework was introduced in our previous works to address the covariance issue in spherically symmetric effective quantum gravity. This paper extends the framework to the electro-vacuum case with a cosmological constant. After analyzing the notion of covariance in the classical theory, we propose an effective Hamiltonian for the electromagnetic field. The effective Hamiltonian together with the effective Hamiltonian constraint of gravity determines an effective dynamical model of gravity coupled to electromagnetic. The resulting model is covariant with respect to both the effective metric and the effective vector potential. By solving the equations of motion derived from the effective Hamiltonian constraint, we obtain several quantum-corrected solutions. Notably, some of these solutions reveal quantum gravity effects manifesting not only in spacetime metrics but also in the electromagnetic field. Finally, the covariance of coupling models with general matter fields is discussed.
@@ I Introduction
As a well-established theory of gravity, general relativity (GR), predicting the formation of black holes (BHs) and describeing the dynamics of the Universe, has successfully passed all experimental and observational tests to date [1, 2]. Nevertheless, the singularities commonly existing according to the singularity theorems [3, 4], due to the geodesic incompleteness of spacetimes, suggest regions where GR breaks down. It is widely expected that a quantum theory of gravity will replace classical GR in these extreme regimes and govern the physics near singularities. As a promising candidate for quantum gravity, loop quantum gravity (LQG) has achieved significant breakthroughs [5, 6, 7, 8, 9, 10, 11, 12]. In particular, both the canonical and covariant formulations of LQG have been individually developed, providing a full understanding of quantum dynamics [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Moreover, the concepts and techniques developed in full LQG have been successfully applied to symmetry-reduced models, giving rise to the fields of loop quantum cosmology (LQC) and loop quantum black holes (LQBH). Within these frameworks, classical singularities are resolved by some natural scenarios, such as big-bounce cosmology and BH-to-white hole transitions [23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Despite these achievements, several issues remain, particularly regarding covariance, which poses a challenge for the effective theory of the quantum symmetry-reduced models. In LQC, covariance is less problematic, as the choice of coordinates adapted to the symmetries of homogeneity and isotropy fully resolves the diffeomorphism constraint. However, in LQBH, while the choice of coordinates adapted to spherical symmetry simplifies the constraint structure, it still leaves the radial component of the diffeomorphism constraint unsolved. Thus, it becomes necessary to consider diffeomorphism transformations along a direction that combines both the temporal and radial coordinates. Consequently, the covariance issue in LQBH models becomes non-trivial and requires careful treatments [32, 33, 34]. This difficulty arises because the effective theory of LQBH is formulated within the Hamiltonian framework, which relies on a 3+1 decomposition of spacetime.
@@
There are several approaches to preserving covariance in effective LQBH models. One approach is to express the equations in terms of tensors, similar to classical GR. An example of this approach is the construction of the quantum Oppenheimer-Snyder model [35, 36]. In this model, a spherically symmetric vacuum is glued to a collapsing ball of spherically symmetric, homogeneous, and pressureless dust. The dynamics of the dust ball is governed by the effective Friedmann equation derived from LQC in [25]. It turns out that the exterior spacetime can be uniquely determined using the Israel junction condition [37, 38], under the assumption that the exterior region remains stationary. Since the Israel junction condition is a tensor equation, the resulting quantum-corrected exterior spacetime remains covariant. In other words, one can choose different coordinate systems in the exterior region during the matching process and still obtain the same exterior spacetime metric tensor. Another approach is to introduce a specific matter field to fix the gauge of diffeomorphism [39, 40]. In this approach, different choices of matter fields for gauge fixing can lead to entirely different expressions for the Hamiltonian constraints. Some of these formulations may fail to incorporate key features, such as polymerization, which characterizes LQG. This raises an important question: what kind of matter fields should be selected to ensure that the associated effective Hamiltonian retains these key LQG features? The third approach focuses on constructing an effective Hamiltonian constraint that preserves covariance, as in the classical theory, as done in [41, 42]. More specifically, the effective Hamiltonian constraint is designed to ensure that the gauge transformations of the metric, generated by the constraints, correspond exactly to its spacetime diffeomorphism transformations. In these works, the approach starts with a modified constraint algebra incorporating quantum gravity effects. The covariance equation for the effective Hamiltonian constraint is then systematically derived based on the necessary and sufficient conditions required to maintain covariance. By solving the covariance equation, one can derive various effective Hamiltonian constraints that correspond to different covariant theories. It is important to note that although these works focus on vacuum gravity, the necessary and sufficient conditions for spacetime covariance are applicable to any matter coupling. Therefore, this approach can be extended to study scenarios beyond the vacuum case, such as the dust-coupling model [41, 42]. The aim of this paper is to demonstrate that the framework developed in [41, 42] can also be extended to the electro-vacuum setting with a cosmological constant. This extension will ensure covariance not only for the spacetime metric but also for the electromagnetic (EM) field.
@@
The remainder of this paper is organized as follows. In Sec. II, we first recall the Hamiltonian formulation of gravity coupled to EM field in the presentation of a cosmological constant, then analyze the notion of covariance in the classical case, and present its exact solutions of the classical model. In Sec. III, we turn to the effective model of gravity coupled to EM field. We construct an effective Hamiltonian for EM field, together with the effective Hamiltonian constraint of gravity, ensuring the covariance of effective model. Several exact solutions of effective model are presented. In Sec. IV, the covariance of coupling models with general matter fields is discussed. We show that the resulting spacetime is always covariant for gravity coupled to an arbitrary matter field in effective theory, once the constraint algebra of coupling model has the similar structure as that of the effective vacuum model. In comparison to gravity, the covariance of matter field needs to be concretely treated. Our results are summarized in Sec. V.
@@ II Classical theory
In this section, we will first recall the Hamiltonian formulation of gravity minimally coupled to EM field in the presentation of a cosmological constant. Building on the Hamiltonian framework, we will check the covariance of the model explicitly. Next, we introduce the reduced formulation of the model with the Gauss constraint solved. Using this reduced formulation, we derive the solutions to the equations of motion (EOM) and examine the covariance of the coupling model based on these solutions.
@@ II.1 Hamiltonian framework of classical theory
Let us focus on the spherically symmetric case. The spacetime is based on a 4-D manifold ×$\mathcal{M}\cong\mathbb{R}\times\Sigma$ with $\Sigma\cong{\mathbb{R}}^{+}\times{\mathbb{S}}^{2}$. Here ${\mathbb{R}}^{+}$ denotes the positive real line, and ${\mathbb{S}}^{2}$ is the 2-sphere. Denote the coordinates of the spacetime by $(t,x,\theta,\phi)$. The phase space consists of the canonical variables $(K_{1},E^{1};K_{2},E^{2})$ for gravity and $(\Gamma,P_{\Gamma})$ for EM field [43, 44, 45, 46, 47]. The nontrivial Poisson brackets between these variables read
@@
with $G$ being the Newtonian constant. The physical meaning of the canonical variables becomes evident from the following facts. The variables $E^{I}$ define the 3-D metric
@@
where the lowercase latin indices $a,b,\dots=1,2,3$ stand for the spatial indices, and ${\rm d}\Omega^{2}={\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}$. The variable $\Gamma$ describes the spatial component of the vector potential of EM field, namely,
@@
The dynamics of the coupling model is governed by a set of constraints: the Hamiltonian constraint $H=H^{GR}+H^{EM}$, the diffeomorphism constraint $H_{x}=H_{x}^{GR}+H_{x}^{EM}$, and the Gauss constraint of EM field $\mathcal{G}=G^{EM}$. Here the superscripts “GR” and “EM” stand for gravity (described by GR) and EM field, respectively. Specifically, we have
@@
$H^{EM}$ $:=\frac{E^{2}P_{\Gamma}^{2}}{2(E^{1})^{3/2}}+\frac{\sqrt{E^{1}}}{E^{2}}\Gamma G^{EM},$ (7)
@@
$G^{EM}$ $:=-\partial_{x}P_{\Gamma},$ (10)
where $\Lambda$ denotes the cosmological constant. These constraints obey the following algebra:
@@
with the structure function $S=\frac{E^{1}}{(E^{2})^{2}}$. Here we introduced the convention $F[g]\equiv\int{\rm d}xF(x)g(x)$ to denote the smeared constraints. It is worth noting that the present paper adopts the EM Hamiltonian $H^{EM}$ in Eq. (7) that differs from the one provided in [46, 47] by a term proportional to the Gauss constraint. This additional term is introduced to ensure that the Hamiltonian constraint and the diffeomorphism constraint form a subalgebra, as demonstrated in Eqs. (11d)–(11f).
@@
Then, the Hamilton’s equations, as the EOM, read
$\dot{K_{1}}$ :=$:=\left\{K_{1},{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}P_{\Gamma}^{2}+N\frac{G}{\sqrt{E^{1}}E^{2}}\Gamma\mathcal{G},$ (13)
$\dot{K_{2}}$ :=$:=\left\{K_{2},{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}P_{\Gamma}^{2}-N\frac{G\sqrt{E^{1}}}{(E^{2})^{2}}\Gamma\mathcal{G},$ (14)
$\dot{E^{1}}$ :=$:=\left\{E^{1},{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (15)
$\dot{E^{2}}$ :=$:=\left\{E^{2},{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\},$ (16)
$\dot{\Gamma}$ :=$:=\left\{\Gamma,{\mathbb{H}}\right\}=\partial_{x}\Psi+\partial_{x}\left[\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\Gamma\right]+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}P_{\Gamma},$ (17)
$\dot{P_{\Gamma}}$ :=$:=\left\{P_{\Gamma},{\mathbb{H}}\right\}=-\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\mathcal{G},$ (18)
@@
stands for the gravitational Hamiltonian. Solving the Hamilton’s equations together with the constraint equations $H=0$, $H_{x}=0$ and $\mathcal{G}=0$, we get the solutions $E^{I}(t,x)$, $K_{I}(t,x)$, $\Gamma(t,x)$, and $P_{\Gamma}(t,x)$ as fields on $\mathcal{M}$. Denote the spacetime indicies by the Greek letters $\rho,\sigma,\dots=0,1,2,3$. Then, the 4-D metric $g_{\rho\sigma}$ and 4-D vector potential $A_{\rho}$ are given by [46, 47]
@@ II.2 Covariance of classical theory
It is observed that $H_{x}[N^{x}]$ generates the diffeomorphism transformation along the vector field $N^{x}\partial_{x}$. Therefore, the Hamilton’s equations (13)–(18) can be rewritten by the Lie derivatives as
@@
where $I=1,2$, $\mathfrak{N}=\partial_{t}-N^{x}\partial_{x}$, and $\mathcal{L}_{\mathfrak{N}}$ denotes the Lie derivative along the vector $\mathfrak{N}$ on $\mathcal{M}$. The fields $K_{1}$, $E^{2}$, $\Gamma$ are scalar densities with weight $1$, and $K_{2}$, $E^{1}$ and $P_{\Gamma}$ are scalars, which is relevant for calculating the Lie derivative. Note that by converting Eqs. (13)–(18) into Eq. (23), we indeed establish the relationship between the evolution of a point in phase space and the corresponding evolution of the fields in spacetime.
Now, let us consider a gauge transformation generated by $H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]$, for arbitrary smeared functions $\alpha$ and $\lambda$, and an arbitrary smeared vector field $\beta^{x}\partial_{x}$. Under the gauge transformation, the phase space variables $W\equiv K_{I},E^{I},\Gamma,P_{\Gamma}$ transform as
where
To show the covariance of the coupling model, we need to demonstrate that the gauge transformations in phase space correspond to the diffeomorphism transformations in spacetime. More precisely, we will prove
@@
Observing that the change $\delta g_{\rho\sigma}$ of $g_{\rho\sigma}$ given by Eq. (20) includes $\delta E^{I}$, $\delta N$ and $\delta N^{x}$, let us first calculate $\delta E^{I}$ by definition (25). Here, as discussed in [41, 42], it is sufficient to consider the case where $\alpha$, $\beta^{x}$ and $\lambda$ are phase space independent. Then, we have
@@
To calculate $\delta N$ and $\delta N^{x}$, we need to consider that the gauge transformation (24) will map a solution $K_{I}(t,x)$, $E^{I}(t,x)$ to a new set of solutions
The new solutions need to satisfy the Hamilton’s equations with respect to a new lapse function $\tilde{N}=N+\epsilon\delta N$ and a new shift vector $\tilde{N}^{x}=N^{x}+\epsilon\delta N^{x}$. This requirement allows us to derive the expressions for $\delta N$ and $\delta N^{x}$, as shown in [41, 42]. Actually, the works [41, 42], which focused on the vacuum case, derived the transformation law for $N$ and $N^{x}$ based on the two key ingredients: (a) the fact that the diffeomorphism constraint generates spatial diffeomorphism transformations, which results in EOM for $E^{I}$ and $K_{I}$ taking the same forms as Eq. (23); and (b) the structure of constraint algebra, which is the same as Eqs. (11d)–(11f). Clearly, these two ingredients (a) and (b) are preserved for the total constraints in the current model. Thus, the transformation laws for $\delta N$ and $\delta N^{x}$ remain the same as in [41, 42], which read
@@
$\delta N^{x}=$ $-N^{2}S\partial_{x}\alpha-(\mathcal{L}_{\beta}\mathfrak{N})^{x}.$
Combing Eqs. (30) and (32), we arrive at Eq. (26).
To demonstrate Eq. (27), where $\delta A_{\rho}$ involves $\delta\Gamma$ and $\delta\Psi$, we begin by analyzing $\delta\Gamma$. Similar to getting Eq. (30), we have
$\tilde{\Gamma}=$ $\Gamma+\epsilon\delta\Gamma+O(\epsilon^{2}),$ (34)
They need to satisfy the Hamilton’s equations with respect to a new lapse function $\tilde{N}=N+\epsilon\delta N$, a new shift vector $\tilde{N}^{x}=N^{x}+\epsilon\delta N^{x}$, and a new Lagrangian multiplier $\tilde{\Psi}=\Psi+\epsilon\delta\Psi$. Notice that $\delta N$ and $\delta N^{x}$ are given by Eq. (32). Then, this requirement will help us to get $\delta\Psi$, with the detailed derivations shown as follows. Substituting $\tilde{\Gamma}$ and $\tilde{P}_{\Gamma}$ given by Eq. (34) into Eq. (23) and comparing the first order results in $\epsilon$ on both sides, we get
+$+\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}},$
Then, we compute the second term on the left-hand side of Eq. (35) and get
$\mathcal{L}_{\mathfrak{N}}\delta X$ (37)
+$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
Note that in the last step in Eq. (37), the first term contains the Lie derivatives for the phase-space dependent parts in $\delta X$, and the second term includes the Lie derivative for the phase-space independent variables. Applying the Jacobi identity for Poisson brackets to Eq. (37), we get
$\mathcal{L}_{\mathfrak{N}}\delta X$ (38)
$=$ $\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}}$
+$+\Big{\{}X,\left\{H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda],H[N]+\mathcal{G}[\Psi]\right\}\Big{\}}$
+$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
The second term on the right-hand side of Eq. (38) contains Poisson bracket between constraints, and thus it can be evaluated by applying the constraint algebra (11a)–(11f). The result reads
$\mathcal{L}_{\mathfrak{N}}\delta X$ (39)
$=$ $\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}}$
+$+\left\{X,\mathcal{G}\left[N\sqrt{S}(-\Psi\partial_{x}\alpha-\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi\right]\right\}$
+$+\Big{\{}X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\Big{\}}.$
Putting Eqs. (36) and (39) into Eq. (35) yields
+$+\Big{\{}X,H[\beta^{x}\partial_{x}N+\mathcal{L}_{\mathfrak{N}}(\alpha N)-\delta N]\Big{\}}$
+$+\left\{X,\mathcal{G}\left[\xi\right]\right\}=0,$
where the coefficient $\xi$ is
$\xi=-N\sqrt{S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)-\delta\Psi.$ (41)
Note that the first two terms in Eq. (40) vanishes due to Eq. (32). Thus, Eq. (40) gives
Thanks to Eqs. (30), (32), (33) and (42), a straightforward computation leads to Eq. (27).

### v3

# thanks: Corresponding author

@@ Abstract
An algebraic framework was introduced in our previous works to address the covariance issue in spherically symmetric effective quantum gravity. This paper extends the framework to the electrovacuum case with a cosmological constant. After analyzing the notion of covariance in the classical theory, we propose an effective Hamiltonian for the electromagnetic field. The effective Hamiltonian together with the effective Hamiltonian constraint of gravity determines an effective dynamical model of gravity coupled to the electromagnetic field. The resulting model is covariant with respect to both the effective metric and the effective vector potential. By solving the equations of motion derived from the effective Hamiltonian constraint, we obtain several quantum-corrected solutions. Notably, some of these solutions reveal quantum gravity effects manifesting not only in spacetime metrics but also in the electromagnetic field. Finally, the covariance of coupling models with general matter fields is discussed.
@@ I Introduction
As a well-established theory of gravity, general relativity (GR), predicting the formation of black holes (BHs) and describing the dynamics of the Universe, has successfully passed all experimental and observational tests to date [1, 2]. Nevertheless, the singularities commonly existing according to the singularity theorems [3, 4], due to the geodesic incompleteness of spacetimes, suggest regions where GR breaks down. It is widely expected that a quantum theory of gravity will replace classical GR in these extreme regimes and govern the physics near singularities. As a promising candidate for quantum gravity, loop quantum gravity (LQG) has achieved significant breakthroughs [5, 6, 7, 8, 9, 10, 11, 12]. In particular, both the canonical and covariant formulations of LQG have been individually developed, providing a full understanding of quantum dynamics [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Moreover, the concepts and techniques developed in full LQG have been successfully applied to symmetry-reduced models, giving rise to the fields of loop quantum cosmology (LQC) and loop quantum BH (LQBH). Within these frameworks, classical singularities are resolved by some natural scenarios, such as big-bounce cosmology and BH-to-white hole transitions [23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Despite these achievements, several issues remain, particularly regarding covariance, which poses a challenge for the effective theory of the quantum symmetry-reduced models. In LQC, covariance is less problematic, as the choice of coordinates adapted to the symmetries of homogeneity and isotropy fully resolves the diffeomorphism constraint. However, in LQBH, while the choice of coordinates adapted to spherical symmetry simplifies the constraint structure, it still leaves the radial component of the diffeomorphism constraint unsolved. Thus, it becomes necessary to consider diffeomorphism transformations along a direction that combines both the temporal and radial coordinates. Consequently, the covariance issue in LQBH models becomes nontrivial and requires careful treatments [32, 33, 34]. This difficulty arises because the effective theory of LQBH is formulated within the Hamiltonian framework, which relies on a $3+1$ decomposition of spacetime.
@@
There are several approaches to preserving covariance in effective LQBH models. One approach is to express the equations in terms of tensors, similar to classical GR. An example of this approach is the construction of the quantum Oppenheimer-Snyder model [35, 36]. In this model, a spherically symmetric vacuum is glued to a collapsing ball of spherically symmetric, homogeneous, and pressureless dust. The dynamics of the dust ball is governed by the effective Friedmann equation derived from LQC in [25]. It turns out that the exterior spacetime can be uniquely determined using the Israel junction condition [37, 38], under the assumption that the exterior region remains stationary. Since the Israel junction condition is a tensor equation, the resulting quantum-corrected exterior spacetime remains covariant. In other words, one can choose different coordinate systems in the exterior region during the matching process and still obtain the same exterior spacetime metric tensor. Another approach is to introduce a specific matter field to fix the gauge of diffeomorphism [39, 40]. In this approach, different choices of matter fields for gauge fixing can lead to entirely different expressions for the Hamiltonian constraints. Some of these formulations may fail to incorporate key features, such as polymerization, which characterizes LQG. This raises an important question: what kind of matter fields should be selected to ensure that the associated effective Hamiltonian retains these key LQG features? The third approach focuses on constructing an effective Hamiltonian constraint that preserves covariance, as in the classical theory, as done in [41, 42, 43, 44]. More specifically, the effective Hamiltonian constraint is designed to ensure that the gauge transformations of the metric, generated by the constraints, correspond exactly to its spacetime diffeomorphism transformations. In these works, the approach starts with a modified constraint algebra incorporating quantum gravity effects. The covariance equation for the effective Hamiltonian constraint is then systematically derived based on the necessary and sufficient conditions required to maintain covariance. By solving the covariance equation, one can derive various effective Hamiltonian constraints that correspond to different covariant theories. It is important to note that although these works focus on vacuum gravity, the necessary and sufficient conditions for spacetime covariance are applicable to any matter coupling. Therefore, this approach can be extended to study scenarios beyond the vacuum case, such as the dust-coupling model [42, 43]. The aim of this paper is to demonstrate that the framework developed in [42, 43] can also be extended to the electrovacuum setting with a cosmological constant. This extension will ensure covariance not only for the spacetime metric but also for the electromagnetic (EM) field.
When including matter coupling, it is crucial, as emphasized in [45], to examine the covariance properties of matter fields—particularly when these fields are tensorial. In [45], the field-strength tensor is chosen as the fundamental variable when imposing the covariance condition, since physical observables must exhibit covariance. While this reasoning is valid, the requirement that physical observables respect covariance can alternatively be implemented by imposing the covariance condition directly on the phase space variables. As we will show, the essential point is that, as gauge fields, these phase space variables are only required to be covariant up to a gauge transformation, which is actually the case in the classical theory. In addition, our approach can be easily generalized to non-Abelian Yang-Mills theory, where the field-strength tensor is no longer gauge invariant.
@@
The remainder of this paper is organized as follows. In Sec. II, we first recall the Hamiltonian formulation of gravity coupled to the EM field in the presentation of a cosmological constant, then analyze the notion of covariance in the classical case, and present its exact solutions of the classical model. In Sec. III, we turn to the effective model of gravity coupled to the EM field. We construct an effective Hamiltonian for the EM field, together with the effective Hamiltonian constraint of gravity, ensuring the covariance of the effective model. Several exact solutions of the effective model are presented. In Sec. IV, the covariance of coupling models with general matter fields is discussed. We show that the resulting spacetime is always covariant for gravity coupled to an arbitrary matter field in effective theory, once the constraint algebra of the coupling model has the similar structure as that of the effective vacuum model. In comparison to gravity, the covariance of matter field needs to be concretely treated. Our results are summarized in Sec. V.
@@ II Classical theory
In this section, we will first recall the Hamiltonian formulation of gravity minimally coupled to the EM field in the presentation of a cosmological constant. Building on the Hamiltonian framework, we will check the covariance of the model explicitly. Next, we introduce the reduced formulation of the model with the Gauss constraint solved. Using this reduced formulation, we derive the solutions to the equations of motion (EOM) and examine the covariance of the coupling model based on these solutions.
@@ II.1 Hamiltonian framework of classical theory
Let us focus on the spherically symmetric case. The spacetime is based on a four-dimensional (4D) manifold $\mathcal{M}\cong\mathbb{R}\times\Sigma$ with $\Sigma\cong\mathcal{M}_{1}\times{\mathbb{S}}^{2}$. Here $\mathcal{M}_{1}$ denotes some one-dimensional manifold which could be, for instance, $\mathbb{R}_{+}$, and ${\mathbb{S}}^{2}$ is the 2-sphere. Denote the coordinates of the spacetime by $(t,x,\theta,\phi)$. The phase space consists of the canonical variables $(K_{1},E^{1};K_{2},E^{2})$ for gravity and $(\Gamma,P_{\Gamma})$ for the EM field [46, 47, 48, 49, 50]. The nontrivial Poisson brackets between these variables read
@@
with $G$ being the Newtonian constant. The physical meaning of the canonical variables becomes evident from the following facts. The variables $E^{I}$ define the 3D metric
@@
where the lowercase latin indices $a,b,\ldots=1,2,3$ stand for the spatial indices and ${\rm d}\Omega^{2}={\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}$. The variable $\Gamma$ describes the spatial component of the vector potential of the EM field, namely,
@@
The dynamics of the coupling model is governed by a set of constraints: the Hamiltonian constraint $H=H^{GR}+H^{EM}$, the diffeomorphism constraint $H_{x}=H_{x}^{GR}+H_{x}^{EM}$, and the Gauss constraint $\mathcal{G}$ of the EM field. Here the superscripts “GR” and “EM” stand for gravity (described by GR) and the EM field, respectively. Specifically, we have
@@
$H^{EM}$ $:=\frac{E^{2}P_{\Gamma}^{2}}{2(E^{1})^{3/2}}+\frac{\sqrt{E^{1}}}{E^{2}}\Gamma\mathcal{G},$ (7)
@@
$\mathcal{G}$ $:=-\partial_{x}P_{\Gamma},$ (10)
where $\Lambda$ denotes the cosmological constant. A straightforward derivation shows that these constraints obey the following algebra:
@@
with the structure function $S=\frac{E^{1}}{(E^{2})^{2}}$. Here we introduced the convention $F[g]\equiv\int{\rm d}xF(x)g(x)$ to denote the smeared constraints. Here, the term proportional to the Gauss constraint of the EM field is included as the diffeomorphism constraint of the EM field, as done in [50, 51]. This term ensures that the diffeomorphism constraint generates spatial diffeomorphism transformation for the EM fields. Moreover, the present paper adopts the EM Hamiltonian $H^{EM}$ in Eq. (7) that differs from the one provided in [49, 50] by a term proportional to the Gauss constraint. This additional term is introduced to ensure that the Hamiltonian constraint and the diffeomorphism constraint form a subalgebra, as demonstrated in Eqs. (11d)–(11f), so that the results in our previous work [42] can be applied directly. It is noted that, as a totally constrained system of the first class, these modifications of constraints do not affect the dynamics.
@@
Then, Hamilton’s equations, as the EOM, read
$\dot{K_{1}}$ $:=\left\{K_{1},{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}P_{\Gamma}^{2}+N\frac{G}{\sqrt{E^{1}}E^{2}}\Gamma\mathcal{G},$ (13)
$\dot{K_{2}}$ $:=\left\{K_{2},{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}P_{\Gamma}^{2}-N\frac{G\sqrt{E^{1}}}{(E^{2})^{2}}\Gamma\mathcal{G},$ (14)
$\dot{E^{1}}$ $:=\left\{E^{1},{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (15)
$\dot{E^{2}}$ $:=\left\{E^{2},{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\},$ (16)
$\dot{\Gamma}$ $:=\left\{\Gamma,{\mathbb{H}}\right\}=\partial_{x}\Psi+\partial_{x}\left[\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\Gamma\right]+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}P_{\Gamma},$ (17)
$\dot{P_{\Gamma}}$ $:=\left\{P_{\Gamma},{\mathbb{H}}\right\}=-\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\mathcal{G},$ (18)
@@
stands for the gravitational Hamiltonian. Solving Hamilton’s equations together with the constraint equations $H=0$, $H_{x}=0$ and $\mathcal{G}=0$, we get the solutions $E^{I}(t,x)$, $K_{I}(t,x)$, $\Gamma(t,x)$, and $P_{\Gamma}(t,x)$ as fields on $\mathcal{M}$. Denote the spacetime indices by the Greek letters $\rho,\sigma,\ldots=0,1,2,3$. Then, the 4D metric $g_{\rho\sigma}$ and the 4D vector potential $A_{\rho}$ are given by [49, 50]
@@ II.2 Covariance of classical theory
It is observed that $H_{x}[N^{x}]$ generates the diffeomorphism transformation along the vector field $N^{x}\partial_{x}$. Therefore, Hamilton’s equations (13)–(18) can be rewritten by the Lie derivatives as
@@
where $I=1,2$, $\mathfrak{N}=\partial_{t}-N^{x}\partial_{x}$, and $\mathcal{L}_{\mathfrak{N}}$ denotes the Lie derivative along the vector $\mathfrak{N}$ on $\mathcal{M}$. The fields $K_{1}$, $E^{2}$, $\Gamma$ are scalar densities with weight $1$, and $K_{2}$, $E^{1}$, and $P_{\Gamma}$ are scalars, which is relevant for calculating the Lie derivative. Note that by converting Eqs. (13)–(18) into Eq. (23), we indeed establish the relationship between the evolution of a point in phase space and the corresponding evolution of the fields in spacetime.
Now, let us consider a gauge transformation generated by $H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]$, for arbitrary smeared functions $\alpha$ and $\lambda$, and an arbitrary smeared vector field $\beta^{x}\partial_{x}$. To show the covariance of the coupling model, we need to demonstrate that the gauge transformations in phase space correspond to the diffeomorphism transformations in spacetime. More precisely, we will prove
@@
Under the gauge transformation, the phase space variables $W\equiv K_{I},E^{I},\Gamma,P_{\Gamma}$ transform as
where $\epsilon$ denotes a small constant that characterizes the parameter of one parameter diffeomorphism generalized by $H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]$, and
As explicitly computed in Appendix A, this yields the following transformation laws:
@@
$\delta E^{2}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}E^{2}-E^{2}\mathfrak{N}^{\rho}\partial_{\rho}\alpha,$
$\delta\Gamma=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}\Gamma+\left(N\frac{\sqrt{E^{1}}}{E^{2}}\Gamma+\Psi\right)\partial_{x}\alpha-\Gamma\mathfrak{N}^{\rho}\partial_{\rho}\alpha+\partial_{x}\lambda.$
Since the expression of the metric $g_{\rho\sigma}$ and the EM field $A_{\rho}$ involves also the Lagrangian multiplier $N$, $N^{x}$, and $\Psi$, which could be phase space independent, a subtle issue arises regarding how these phase-space-independent variables transform under gauge transformations. This issue has been thoroughly discussed in [42, 43], where we show that the transformation laws of the phase-space-independent variables are determined by the requirement that $\tilde{W}$, defined in Eq. (26), satisfies the equation of motion (23) associated with the transformed Lagrange multipliers $N+\epsilon\delta N$, $N^{x}+\epsilon\delta N^{x}$, and $\Psi+\epsilon\delta\Psi$. As explicitly derived in Appendix A, this requirement leads to the following transformation rules:
@@
$\delta N^{x}=$ $-N^{2}S\partial_{x}\alpha-(\mathcal{L}_{\beta}\mathfrak{N})^{x},$
$\delta\Psi=$ $-N\sqrt{S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda).$
Combining Eqs. (28) and (29), we can directly show (24) and (25).